OpenStudy (anonymous):

In 1989, Carl Lewis established a world record when he ran the 100. meter dash in 9.92 seconds. What was his average speed (in m/s) for the race? Remember to include your data, equation, and work when solving this problem.

OpenStudy (anonymous):

Vavg=s/t=100m/9.92s=10.08m/s